Tom is 5 times as old as Todd. In 8 years Tom's age will be 22 years more than 2 times as old as Todd. How old is Todd?
Thank you very much for you help

Answer by scott8148(6628)

let x=Todd, so 5x=Tom

"In 8 years Tom's age will be 22 years more than 2 times as old as Todd"
__ 5x+8=2(x+8)+22

5x+8=2x+16+22 __ 3x=30